OpenStudy (anonymous):

a sphere has a 3m diameter. what is its volume to the nearest hundredth?

Directrix (directrix):

The formula for the volume of a sphere is: V = (4/3) * pi * r^3 where r is the radius of the sphere. @carolinagrl

OpenStudy (anonymous):

use the volume formula!

OpenStudy (anonymous):

Directrix has put the formula for you just put the value of (r)

Directrix (directrix):

The radius is 3/2 because the radius is one-half the diameter. V = (4/3) * pi * r^3 V = (4/3) * pi * (3/2)^3 Crank this out and let us know what you get.

OpenStudy (anonymous):

where did you get the 4/3?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

>>>where did you get the 4/3? It is part of the formula for the volume of a sphere.

Directrix (directrix):

Click on attachment to see formula.

OpenStudy (anonymous):

ok

Directrix (directrix):

There is an online calculator for the volume of a sphere given its radius. Look here if you want to use it. https://www.google.com/?gws_rd=ssl#q=volume+of+a+sphere

OpenStudy (anonymous):

0.52 is the answer